Main INDEX
Monthly INDEX
PREV
NEXT
User name Ramesh
Log entry time 23:22:10 on June12,2005
Entry number 145972
keyword=Y_target puzzle of right hrs solved
Now we can see very good y_target spectrum of righ hrs (fig 1), figure 2 is the same for the left hrs.
Right arm Y-target spectrum was not nice since the run 4473 through 4552 . This was because the right hrs sieve was partially in. It was moved out of the beam path in a controlled access just a minute ago. The plots below are from the run 4553 after the controlled access.
FIGURE 1
FIGURE 2